OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 4, –2, –8, … ?

OpenStudy (anonymous):

ummm.?

OpenStudy (anonymous):

Answer –122 –128 –110 –116 those are my choices with ur number im not getting any of them

OpenStudy (anonymous):

4 - (20 - 1) * 6 = -110 .

hero (hero):

That is correct

OpenStudy (anonymous):

-110

OpenStudy (anonymous):

can someone explain how its done?

hero (hero):

Yeah, tnsgod, what's the general formula?

hero (hero):

Okay, I think I got it: Start with the first term, subtract (number for terms minus 1) multiply it by the distance between the terms

OpenStudy (anonymous):

Well, 4 is the first term and each subsequent term is 6 less than the one before. Therefore: ith Term = 4 - (i - 1) * 6

OpenStudy (anonymous):

if you subtract 6 from 4 20 times.. you'll have -110. 4-6 = -2 -2-6 = -8 -8-6 = -14 etc.